OpenStudy (anonymous):

Simultaneous equation 0=C1+C2+C3 0=C1+2C2+3C3 2=C1+4C2+9C3

OpenStudy (anonymous):

solve Q1 for C1 then put the of C1 in both Q2 and Q3

OpenStudy (anonymous):

C1=-(C2+C3) Q2 -C1=2C2+3C3 *** C2+C3=2C2+3C3 Q3 2=C1+4C2+9C3 2-C1=4C2+9C3 *** 2+C2+C3=4C2+9C3